# Iterated sampling importance resampling with adaptive number of proposals

### v1

@@ Appendix H Some properties on Hilbert spaces and reversible Markov chains
Thus it holds that $\langle f|A_{\mathbb{C}}f\rangle_{\mathbb{C}}\geq 0$ for all $f\in\mathcal{H}_{\mathbb{C}}$ for which the earlier result holds and therefore for all $f\in\mathcal{H}$, $\|f\|>0$
@@
Result holds for complex Hilbert space [e.g. rudin, Theorem 12.12]. In the case of Real Hilbert space $\mathcal{H}$ let us do complexification as in proof of Lemma 54. Then $\langle f|A_{\mathbb{C}}f\rangle\geq 0$ for all $f\in\mathcal{H}_{\mathbb{C}}$ and by Lemma 53 $A_{\mathbb{C}}$ is self-adjoint. Additionally for $f,g\in\mathcal{H}$
@@
where the equality holds for $g=A^{-1}f$. By Lemma 54 equation (22) is equal to zero if and only if $\|A^{-1}f-g\|=0$. ∎
@@
Let $J\subset\mathbb{R}$ and $(A_{\lambda})_{\lambda\in J}$ be collection of bounded self-adjoint operators on Hilbert space $\mathcal{H}$ for which $\langle f|A_{\lambda}f\rangle\geq 0$ for all $f\in H$, the inverse $A_{\lambda}^{-1}$ exists and $\lambda\mapsto\langle f|A_{\lambda}f\rangle$ is injective for $\|f\|>0$, when $\lambda\in J$.
@@
Follows from Lemma 56 and 57. ∎
@@
    $u_{g}$ is decreasing (increasing) function for all $g\in L_{0}^{2}(\pi)$.
@@
    $u_{g}$ is convex function for all $g\in L_{0}^{2}(\pi)$, where $J$ is interval.
@@
    $v_{f}$ is decreasing (increasing) function if (a) holds,
@@
    $v_{f}$ is strictly decreasing (increasing) function if (b) holds,
@@
    $v_{f}$ is convex function if (c) holds,
@@
    $v_{f}$ is strictly convex function if (d) holds,
@@
Since $Q_{\lambda}$ is $\pi$-reversible we have $I-Q_{\lambda}:L_{0}^{2}(\pi)\to L_{0}^{2}(\pi)$ is self-adjoint. For $g\in L_{0}^{2}(\pi)$ let us denote the Dirichlet form by
@@
which is non-negative for the Markov kernel and the function $\lambda\mapsto-\mathcal{E}_{\lambda}(g)$ preserves conditions in (a)–(d). Let us denote $h_{\lambda}=(I-Q_{\lambda})^{-1}f$, for which holds by Lemma 56
@@
Thus for $\in J, x<yx<y$, when (a) holds with the decreasing case (for $y<x$, when (a) holds with the increasing case)
@@
where the last inequality is strict if (b) holds. For $\in J, x<yx<y$ and $t\in(0,1)$, when (c) holds
@@
where the first inequality is strict if (d) holds. However if (b) and (c) hold, then also (A3) holds and

### v2

@@ Appendix H Some properties on Hilbert spaces and reversible Markov chains
Thus, it holds that $\langle f|A_{\mathbb{C}}f\rangle_{\mathbb{C}}\geq 0$ for all $f\in\mathcal{H}_{\mathbb{C}}$ for which the earlier result holds and therefore for all $f\in\mathcal{H}$, $\|f\|>0$
@@
Result holds for complex Hilbert space [e.g. rudin, Theorem 12.12]. In the case of Real Hilbert space $\mathcal{H}$ let us do complexification as in proof of Lemma 58. Then $\langle f|A_{\mathbb{C}}f\rangle\geq 0$ for all $f\in\mathcal{H}_{\mathbb{C}}$, and by Lemma 57 $A_{\mathbb{C}}$ is self-adjoint. Additionally for $f,g\in\mathcal{H}$
@@
where the equality holds for $g=A^{-1}f$. By Lemma 58, equation (22) is equal to zero if and only if $\|A^{-1}f-g\|=0$. ∎
@@
Let $J\subset\mathbb{R}$ and $(A_{\lambda})_{\lambda\in J}$ be collection of bounded self-adjoint operators on Hilbert space $\mathcal{H}$ for which $\langle f|A_{\lambda}f\rangle\geq 0$ for all $f\in H$, the inverse $A_{\lambda}^{-1}$ exists and $\lambda\mapsto\langle f|A_{\lambda}f\rangle$ is injective for $\|f\|>0$ when $\lambda\in J$.
@@
Follows from Lemma 60 and 61. ∎
@@
    $u_{g}$ is decreasing (increasing) function for all $g\in L_{0,+}^{2}(\pi)$.
@@
    $u_{g}$ is convex function for all $g\in L_{0,+}^{2}(\pi)$, where $J$ is interval.
@@
    $v_{f}$ is decreasing (increasing) function if (a) holds.
@@
    $v_{f}$ is strictly decreasing (increasing) function if (b) holds.
@@
    $v_{f}$ is convex function if (c) holds.
@@
    $v_{f}$ is strictly convex function if (d) holds.
@@
Since $Q_{\lambda}$ is $\pi$-reversible, we have $I-Q_{\lambda}:L_{0}^{2}(\pi)\to L_{0}^{2}(\pi)$ is self-adjoint. For $g\in L_{0}^{2}(\pi)$, let us denote the Dirichlet form by
@@
which is non-negative for the Markov kernel, and the function $\lambda\mapsto-\mathcal{E}_{\lambda}(g)$ preserves conditions in (a)–(d). Let us denote $h_{\lambda}=(I-Q_{\lambda})^{-1}f$. By Lemma 60
@@
Let $x,y\in J$. When (a) holds with the decreasing case, and $x<y$ (or (a) holds with the increasing case, and $x>y$), then
@@
where the last inequality is strict if (b) holds. When (c) holds, $x<y$ and $t\in(0,1)$, then
@@
where the first inequality is strict if (d) holds. However, if (b) and (c) hold, then also (A3) holds and